# Large Language Models are Good Spontaneous Multilingual Learners: Is the Multilingual Annotated Data Necessary?

### v1

@@ Abstract
Recently, Large Language Models (LLMs) have shown impressive language capabilities. However, most of the existing LLMs are all English-centric, which have very unstable and unbalanced performance across different languages. Multilingual alignment is an effective method to enhance the LLMs’ multilingual capabilities. In this work, we explore the multilingual alignment paradigm which utilizes translation data and comprehensively investigate the spontaneous multilingual improvement of LLMs. We find that LLMs only instruction-tuned on question translation data without annotated answers are able to get significant multilingual performance enhancement even across a wide range of languages unseen during instruction-tuning. Additionally, we utilize different settings and mechanistic interpretability methods to comprehensively analyze the LLM’s performance in the multilingual scenario. Our code and data is available at: https://github.com/Shimao-Zhang/LLM-Multilingual-Learner.
Large Language Models are Good Spontaneous Multilingual Learners:
Is the Multilingual Annotated Data Necessary?
@@ 1 Introduction
Large Language Models (LLMs) have recently shown impressive language capabilities across numerous downstream language tasks (Zhao et al., 2023). However, most existing LLMs are trained on extensive high-resource languages text such as English, Chinese, German, French, and so on (Touvron et al., 2023; Brown et al., 2020; Jiang et al., 2023), which lead to a significant performance gap between high-resource languages and low-resource languages (Huang et al., 2023; Zhang et al., 2023b; Gao et al., 2024). For the same task and question contents, using different languages for inputs may have a significant impact on the model’s performance.
@@
Some studies have conducted comprehensive exploration about how to enhance the capabilities of LLMs for low-resource language. The classical approach typically follows the translate-based paradigm (Liu et al., 2024), which translating non-English inputs into English or translating English data into non-English for instruction tuning. However, it is difficult to accurately translate all texts into the target low-resource language (Zhu et al., 2024), not to mention the increasing translation cost as the data scale expands. In order to enhance low-resource languages performance with less cost, some cross-lingual alignment and transfer strategies have been proposed (Eronen et al., 2023; Zhu et al., 2024; Zhao et al., 2024a). But all these methods rely on the data in the target language.
@@
Meanwhile, some studies have explored English-centric LLMs, revealing that English also participate in the intermediate latent reasoning process of these models even when LLMs are prompted in non-English (Wendler et al., 2024; Zhao et al., 2024b). These findings suggest that for LLMs, multiple languages are not completely isolate, and LLMs have the capability to leverage the connections between various languages to address problems in multilingual scenarios. It further demonstrates the feasibility of cross-lingual capability transfer. More surprisingly, Kew et al. (2023) discover that when instruction-tuning LLMs with multilingual data, it is not necessary to instruction-tune the model on data from all target languages to achieve a multilingual ability similar to models instruction-tuned on all languages.
Intuitively, LLMs have abilities and motivations to acclimatize themselves to multilingual environment (Shi et al., 2022). But what should we do to facilitate LLMs to do this better? Most existing methods rely on instruction-tuning on the corresponding datasets (Kew et al., 2023; Liu et al., 2024). However, given the strong capabilities of models in high-resource languages, which indicates that LLMs actually possesses the ability and knowledge to solve problems, such extensive additional annotated data might not be necessary to help LLMs improve their multilingual abilities.
@@
In this work, we further investigate the multilingual learning capabilities of LLMs, where we only train the LLMs on the parallel data without annotated answers (only queries) in a few languages. Following this pattern, we conduct the experiments on models in different types (English-centric or not) and parameter sizes, and test their multilingual capability across a wide range of languages on different benchmarks. We find that multilingual question alignment following Zhu et al. (2024) can effectively enhance the multilingual capabilities of LLMs. Our results indicate that only tuning on questions (without annotated answers) in a small number of languages can bring significant multilingual improvements even across many languages unseen during instruction-tuning process, which implies good language generalization. Furthermore, we also use logit lens (Nostalgebraist, 2020) and dimensionality reduction (Pearson, 1901; Hotelling, 1933) techniques to study the latent states of LLMs, providing more comprehensive perspectives and empirical results for understanding the multilingual learning of large language models.
@@ 2.1 Unbalanced Multilingual Performance
With a much larger number of parameters pretrained on a massive corpus, LLMs have shown the impressive capabilities in a variety of language tasks (Zhao et al., 2023). These models are mainly pretrained on English data, which often accounts for 90% or even more of all training data, such as LLaMA2 (Touvron et al., 2023), GPT-3 (Brown et al., 2020), Mistral (Jiang et al., 2023), Falcon (Almazrouei et al., 2023), and so on. We present a partial language distribution of LLaMA-2’s training data in Table 7 in Appendix A. Meanwhile, most of the LLMs also show unstable and unbalanced performance in multilingual scenarios, especially for some low-resource languages (Zhang et al., 2023a; Zhu et al., 2024). It’s important to enable LLMs to adapt to a wider range of users and scenarios.
@@ 2.2 Cross-lingual Enhancement for Large Language Models
However, LLMs still exhibit significant shortcomings in multilingual scenarios. Many researchers propose multilingual LLMs that are specifically adjusted for multilingual tasks (Team, 2023; Le Scao et al., 2023; Wei et al., 2023). But for multilingual LLMs, researches indicate a decline in their performance in English because of the limited tokens and parameter size (Lin et al., 2022; Scao et al., 2022).
Based on the existing LLMs, researchers have made great efforts to enhancing the multilingual performance of LLMs, which include two categories: prompting close-source LLMs and instruction-tuning open-source LLMs. For the former, some studies utilize translation-based strategies which prompt ChatGPT to translate the non-English input into English firstly before solving the problem (Huang et al., 2023; Qin et al., 2023). This type of approaches are constrained by the translation quality of the model itself and is cumbersome for users.
For the latter, LLMs shows significant improvement of multilingual and multitask abilities and good task generalization through multilingual multitask fine-tuning (Muennighoff et al., 2022). Chen et al. (2023) follow the translation-based approach and instruction-tune the model on a multilingual version of GSM8K, which is translated from English GSM8K (Cobbe et al., 2021). Liang et al. (2024) create a new intermediate language MUL (Machine-created Universal Language) as a translatable unified representation of shared concepts across different languages. "X-English" parallel translation data have also been widely used (Zhu et al., 2024). In our work, we mainly use this type of data, i.e. translation data between two different languages, to enhance multilingual alignment.
@@ 4 Experimental Setup
    Obvious Answers: Obvious answers reduce the entropy during inference process, minimizing the impact of irrelevant tokens on our analysis.
@@
Based on these, we conduct our experiments on two types of tasks:
@@
We all use LoRA (Hu et al., 2021) to instruction-tune the pre-trained models on mixed parallel data first. We train LLMs on the translation data excluding the golden answers to mitigate the impact of the data of the tasks themselves on the model’s capabilities. For controlling the output more flexibly and the reproducibility, we use in-context learning rather than fine-tuning. We use constrained decoding for generation to eliminate the interference of irrelevant outputs on the results. Considering we mainly focus on the language understanding and task solving capabilities, we use English outputs uniformly if it is not specified.
More details are shown in Appendix B.
@@
Qwen1.5-14B en zh de fr es it nl ja ru sv
base 84.50 83.50 74.17 75.17 81.17 78.67 78.17 51.17 76.83 76.17
zh/de $\Rightarrow$ en 91.83 84.50 83.67 84.50 85.00 83.67 84.5 57.17 81.67 84.33
Qwen1.5-14B sl pl bg no ms is hi th sw bn
base 63.17 67.83 64.33 43.00 75.00 48.17 61.00 69.67 45.83 41.33
@@ 5 Results
zh/es $\Rightarrow$ it 73.22 86.38
zh/es $\Rightarrow$ en 76.13 90.83
Table 4: Accuracy on Amazon Reviews Polarity. We replace English with Italian as the target language. Mistral is English-centric and Qwen1.5 is not English-centric.
@@ 5.1 Main Results
    High-resource languages are not only good learners but also good leaders. Is there any difference when we use high-resource languages or low-resource languages in our training data? Our results in Table 1 show that three models trained on Swahili data achieve the top three highest accuracy on Swahili, while the accuracy on high-resource language is not significantly related to whether the corresponding language data is used. More importantly, training on high-resource language data enables the model to achieve more stable improvements across multiple languages compared to that on low-resource languages.
@@
    A few languages are enough for spontaneous multilingual learning. We select one, two, three languages with English for instruction-tuning respectively. As seen in Table 1, although using more languages sometimes leads to more stable improvements, model trained only on Chinese and English have achieved similar overall performance improvements. This is also consistent with the findings of Kew et al. (2023).
@@
    Our findings remain consistent across language models of different parameter sizes. We also present the average accuracy results of Qwen1.5-1.8B, Qwen1.5-4B, and Qwen1.5-14B in Table 2 to strengthen our conclusions. We find significant multilingual performance improvements across all of these models.
We have also validated our findings on the other task, Natural Language Inference (NLI). In this task, the model needs to determine the relationship between the hypothesis and the premise as entailment, neutral, or contradiction. We conduct our experiment on SNLI and report the accuracy of Qwen1.5-14B on natural language inference task in Table 3. Similar to the emotion classification task, we can see that models instruction-tuned on multilingual translation data significantly outperform the base model across these languages, which confirms that our findings have good generalization across different tasks.
@@ 5.2 Analysis
As elaborated in Section 4, we use outputs in English uniformly for all languages in our previous experiments. English has been widely used for multilingual transfer as a pivot language (Zhu et al., 2024; Hu et al., 2023). We further investigate the case of replacing English with Italian and report the results in Table 4. Mistral is an English-centric LLM and Qwen1.5 is not an English-centric LLM. From the results, we can find that using Italian as target language leads to different performances on different types of models. For English-centric LLM, using non-English language as target language has a negative impact on the overall multilingual capabilities of the model. On the contrary, using Italian rather than English is also helpful for Qwen’s multilingual performance improvement, though worse than using English beceuse of the model’s worse capability of Italian than English.
@@
(a) Chinese Before
(b) Japanese Before
(c) Russian Before
(d) Chinese After
(e) Japanese After
(f) Russian After
Figure 1: Logit lens on Mistral-7B in Chinese, Japanese and Russian scenarios. The horizontal axes is the layer num and the vertical axes is the probability. "en" (Orange) means the latent English output corresponding to the correct answer in the target language. "zh"/"ja"/"ru" (Blue) means the correct answer in the target language. "all_possible_out" (Cyan) means the probability of all possible outputs in the target language. "all_possible_latent" (Gray) means the probability of all possible outputs in English.
@@ 6.1 Logit Lens
Following Wendler et al. (2024), we utilize logit lens to analyze the changes of the model. We utilize logit lens on Qwen1.5, a series of LLMs that are not English-centric, and find there is not English latent outputs in the intermediate layers. And the prefix token overlapping between target language and English will also bring errors to the results. So we choose Chinese, Japanese and Russian as three representative languages for our experiment, which shows significant improvement in our results before. Following Wendler et al. (2024), we use the outputs in the same language with the inputs (results shown in Table 6). We conduct our experiments on Mistral-7B and its best trained version "sw/hi $\Rightarrow$ en" in Table 6. We report the results in Figure 1. Clearly, we can observe the following points: (1) All models generate latent English output before generating outputs in the target language finally; (2) The proportion of the probability of the correct answer increases in the sum of all possible answer probabilities; (3) The probability of all other possible answers (except correct answer) in the latent English outputs is nearly zero; (4) The area of latent English output significantly increases, which means the trained models perform inference in English better.
@@ 6.2 Principal Component Analysis
We further utilize the dimensionality reduction technique to visualize the intermediate layer latent outputs of the model across different languages. PCA (Principal Component Analysis) is a normal linear dimensionality reduction technique (Pearson, 1901; Hotelling, 1933), which can be used in some scenarios where logit lens doesn’t work. Principal components are a few linear combinations of the original variables that maximally explain the variance of all the variables (Greenacre et al., 2022). We utilize PCA to map the latent logits into the two-dimensional representation. Based on the patterns shown in Figure 1, we report layer 20, layer 25, layer 30 and the last layer as four representative layers in Figure 2. We have the following findings: (1) The points of different languages follow the similar patterns in layer 20 and layer 25, where English latent outputs have appeared and outputs in the target language haven’t appeared. We further calculate the Pearson correlation coefficient of 1 dimension PCA results (Appendix C). There is a strong linear correlation between representations of different languages, which also indicates an uniform latent representation pattern during inference process; (2) Representations belong to different languages exhibit greater differences from each other in the trained model; (3) The results of the last layer is similar because of the same possible outputs; (4) Based on Pearson coefficient reported in Appendix C, the correlation between Hindi (low-resource language) and other languages (high-resource language) significantly improves.

### v2

# Getting More from Less: Large Language Models are Good Spontaneous Multilingual Learners

@@ Abstract
Recently, Large Language Models (LLMs) have shown impressive language capabilities. While most of the existing LLMs have very unbalanced performance across different languages, multilingual alignment based on translation parallel data is an effective method to enhance the LLMs’ multilingual capabilities. In this work, we discover and comprehensively investigate the spontaneous multilingual alignment improvement of LLMs. We find that LLMs instruction-tuned on the question translation data (i.e. without annotated answers) are able to encourage the alignment between English and a wide range of languages, even including those unseen during instruction-tuning. Additionally, we utilize different settings and mechanistic interpretability methods to analyze the LLM’s performance in the multilingual scenario comprehensively. Our work suggests that LLMs have enormous potential for improving multilingual alignment efficiently with great language and task generalization.  [^1] Our code and data is available at: https://github.com/Shimao-Zhang/LLM-Multilingual-Learner.
Getting More from Less: Large Language Models are Good
Spontaneous Multilingual Learners
@@ 1 Introduction
Large Language Models (LLMs) have recently shown impressive language capabilities across numerous downstream language tasks (Zhao et al., 2023). However, most existing LLMs are trained on extensive high-resource languages text (Touvron et al., 2023; Brown et al., 2020; Jiang et al., 2023), which lead to a significant performance gap between high-resource languages and low-resource languages (Huang et al., 2023; Zhang et al., 2023b; Gao et al., 2024). For the same task and question contents, using different languages for inputs may have a significant impact on the model’s performance.
@@
Some studies have conducted comprehensive exploration about how to enhance the LLMs’ capabilities across different languages. The classical approach typically follows the translate-based paradigm (Liu et al., 2024). Considering LLMs’ great performance on the high-resource languages, some cross-lingual alignment and transfer methods are proposed (Eronen et al., 2023; Zhu et al., 2024; Zhao et al., 2024a). Question alignment (Zhu et al., 2024) is an outstanding paradigm among these methods which effectively improves multilingual alignment at lower cost, i.e. only utilizes the X-English parallel question translation data.
@@
Meanwhile, some studies have further explored the LLMs, revealing that English also participate in the intermediate latent reasoning of these models even when LLMs are prompted in non-English (Wendler et al., 2024; Zhao et al., 2024b). These findings suggest that for LLMs, different languages are not isolated, and LLMs are able to leverage the connections between various languages to address problems in the multilingual scenarios. Researchers also reveal the shared semantic space for different languages (Chang et al., 2022), which is consistent with the findings above and indicates the importance of the multilingual alignment. Also, Kew et al. (2023) discover that multilingual instruction-tuning with three languages improves model’s cross-lingual transfer abilities on some generative tasks.
Intuitively, LLMs have abilities to acclimatize themselves to the multilingual environment through appropriate training (Shi et al., 2022). Many existing methods rely on instruction-tuning on the multilingual instruction-tuning datasets (Kew et al., 2023; Liu et al., 2024). However, given the question alignment paradigm, utilizing multilingual alignment is also helpful for improving LLMs’ multilingual abilities. Additionally, we focus on question alignment in our work to eliminate the interference of task-related data with annotated answers from our analysis of multilingual alignment. Based on the findings above, can LLMs achieve better multilingual alignment across different languages efficiently through appropriate methods?
@@
In this work, we investigate the multilingual alignment of LLMs, where we only train the LLMs on the parallel data without annotated answers (only queries) in a few languages. Following question alignment, we conduct the experiments on models in different types (English-centric or not) and parameter sizes, and test across a wide range of languages on different benchmarks. We find that question alignment following Zhu et al. (2024) can effectively enhance the multilingual capabilities of LLMs, which indicates that models can effectively utilize the relevant knowledge and capabilities learned during the pretraining process with question alignment, consisting with the "Superficial Alignment Hypothesis" (Zhou et al., 2024). Our results also indicate that conducting question alignment in a small number of languages brings significantly better multilingual alignment even between English and many languages unseen during instruction-tuning process, which implies good language generalization. Furthermore, we also use logit lens (Nostalgebraist, 2020) and dimensionality reduction techniques (Pearson, 1901) to study the latent states of LLMs, providing more comprehensive perspectives and empirical results for the alignment improvements in our experiments.
@@ 2.1 Unbalanced Multilingual Performance
With a much larger number of parameters pretrained on a massive corpus, LLMs have shown the impressive capabilities in a variety of language tasks (Zhao et al., 2023). These models are mainly pretrained on English data, which often accounts for 90% or even more of all training data. We present a partial language distribution of LLaMA-2’s training data in Table 7 in Appendix A. Meanwhile, most of the LLMs also show unstable and unbalanced performance in multilingual scenarios, especially for some low-resource languages (Zhang et al., 2023a; Zhu et al., 2024). It’s important to enable LLMs to adapt to a wider range of users and scenarios.
@@ 2.2 Cross-lingual Enhancement for Large Language Models
While LLMs still exhibit significant shortcomings in multilingual scenarios, many researchers propose multilingual LLMs that are specifically adjusted for multilingual tasks (Team, 2023; Le Scao et al., 2023; Wei et al., 2023). But for multilingual LLMs, researches indicate a decline in their performance in English because of the limited tokens and parameter size (Lin et al., 2022; Scao et al., 2022).
Based on the existing LLMs, researchers have made great efforts to enhancing the multilingual performance, which include two categories: prompting close-source LLMs and instruction-tuning open-source LLMs. For the former, some studies utilize translation-based strategies which translate the non-English input into English firstly before solving the problem (Huang et al., 2023; Qin et al., 2023). This type of approaches are constrained by the translation quality of the model itself and is cumbersome for users.
For the latter, LLMs shows significant improvement of multilingual abilities and good task generalization through multilingual multitask fine-tuning (Muennighoff et al., 2022). Chen et al. (2023) follow the translation-based approach and instruction-tune the model on a multilingual version of GSM8K, which is translated from English GSM8K (Cobbe et al., 2021). Liang et al. (2024) create a new intermediate language MUL (Machine-created Universal Language) as a translatable unified representation of shared concepts across different languages. "X-English" parallel question translation data have also been used for multilingual question alignment (Zhu et al., 2024). In our work, we mainly analyse based on the question alignment, which is an outstanding alignment methods, and eliminates the interference of the annotated answers from our analysis.
@@ 4 Experimental Setup
    Obvious Answers: Obvious answers reduce the entropy during the inference process, minimizing the impact of irrelevant tokens on our analysis.
@@
Based on these, we conduct our experiments on three different types of tasks:
@@
  • Paraphrase Identification: Model needs to judge if two given sentences are semantically equivalent in the paraphrase identification task, which includes two possible labels. We conduct our experiments on PAWS [^4] https://huggingface.co/datasets/google-research-datasets/paws dataset (Zhang et al., 2019), which is a famous dataset proposed by Google. Following the above tasks, we extract 10K instances from train subset for parallel data and 500 instances from test subset for test data respectively.
@@
We all use LoRA (Hu et al., 2021) to instruction-tune the pre-trained models on the mixed parallel translation data first. We train LLMs on the translation data excluding the golden answers to mitigate the impact of the data of the tasks themselves on the model’s capabilities. We use in-context learning which not only doesn’t interfere with LLMs’ parameters but also help LLMs handle the tasks better. We use constrained decoding rather than sampling that is used for diverse generation (Zhang et al., 2024) to eliminate the interference of irrelevant outputs on the results. More details are shown in Appendix B.
@@
Qwen1.5-14B Emotion Classification NLI Paraphrase Identification
base 86.27 66.94 76.36
zh/es $\Rightarrow$ en 91.53 73.03 82.59
zh/de $\Rightarrow$ en 92.25 73.28 82.21
zh/it $\Rightarrow$ en 92.13 71.94 82.15
sw/hi $\Rightarrow$ en 90.28 71.48 81.80
Table 3: Average accuracy results of Qwen1.5-14B base model and trained models on the Amazon Reviews Polarity, SNLI and PAWS across 20 different languages. The accuracy of randomly choosing is 33.33% for SNLI and 50.00% for the other two tasks. We highlight the best results for every task. Full results are reported in Appendix E.
@@ 5 Results
ja/it $\Rightarrow$ zh 73.32 90.34
sw/hi $\Rightarrow$ zh 77.00 90.58
en/ja $\Rightarrow$ zh 71.92 90.04
Table 4: Average accuracy on Amazon Reviews Polarity. We replace English with Chinese as the target language. We highlight the best results for each model.
@@ 5.1 Main Results
    High-resource languages are not only good learners but also good leaders. Is there any difference when we use high-resource languages or low-resource languages in our training data? Our results in Table 1 show that the accuracy on high-resource language is not significantly related to whether the corresponding language data is used. More importantly, training on high-resource language data enables the model to achieve more stable improvements across multiple languages compared to that on low-resource languages (Swahili, Hindi, and Thai).
@@
    A few languages can lead to spontaneous multilingual learning. We select one, two, three languages with English for instruction-tuning respectively. In Table 1, although using more languages sometimes leads to more stable improvements, model trained only on Chinese and English have achieved similar overall performance improvements. This is also consistent with the findings of Kew et al. (2023). The multilingual alignment improvement shows great language generalization.
@@
    Our findings remain consistent across models of different parameter sizes. We also present the average accuracy results of Qwen1.5-1.8B, Qwen1.5-4B, and Qwen1.5-14B in Table 2 to strengthen our conclusions. We find significant multilingual performance improvements across all of these models.
We have also validated our findings on the other two tasks to strengthen our conclusions, including Natural Language Inference (NLI) and Paraphrase Identification. The model needs to determine the relationship between two paragraphs of text in both of these two tasks. We conduct our experiment on SNLI for NLI task and PAWS for Paraphrase Identification task. We report the average accuracy of Qwen1.5-14B across all languages in Table 3. And we report the full results on each language in Appendix E.
@@ 5.2 Analysis
As elaborated in Section 4, we use outputs in English uniformly for all languages in our previous experiments. English has been widely used for multilingual transfer as a pivot language (Zhu et al., 2024; Hu et al., 2023). We further investigate the case of replacing English with Chinese in training data and report the results in Table 4. Mistral and Qwen1.5 represent two different types of LLMs (English-centric or not) respectively. From the results, we can find that using Chinese as the target language leads to the same conclusions as using English. For both of the two types of LLMs, using Chinese rather than English as the target language is also helpful for models’ multilingual performance improvement, which indicates that English is not necessary as the target language in the training data.
(a) Chinese Before
(b) Japanese Before
(c) Russian Before
(d) Chinese After
(e) Japanese After
(f) Russian After
Figure 1: Logit lens on Mistral-7B in Chinese, Japanese and Russian scenarios (languages not in training data). The horizontal axes is the layer num and the vertical axes is the probability. "en" (Green covered by Red) means the latent English output corresponding to the correct answer in the target language. "zh/ja/ru" (Orange) means the correct answer in the target language. "all_possible_out" (Blue) means the probability of all possible outputs in the target language. "all_possible_latent" (Red) means all possible outputs in English.
@@ 6.1 Logit Lens
Following Wendler et al. (2024), we utilize logit lens to analyze the changes of the model. We utilize logit lens on Qwen1.5, a series of LLMs that are not English-centric, and find there is not English latent outputs in the intermediate layers. And the prefix token overlapping between target language and English will also bring errors to the results. So we choose Chinese, Japanese and Russian as three representative languages for our experiment, which shows significant improvement in our results before. Following Wendler et al. (2024), we use the outputs in the same language with the inputs (Table 6). We conduct our experiments on Mistral-7B and its best trained version "sw/hi $\Rightarrow$ en" in Table 6. We report the results in Figure 1. Clearly, we can observe the following points: (1) All models generate latent English output before generating outputs in the target language finally; (2) The proportion of the probability of the correct answer increases in the sum of all possible answer probabilities; (3) The probability of all other possible answers (except correct answer) in the latent English outputs is nearly zero; (4) The area of latent English output significantly increases, which means the trained models perform latent inference in English better and indicates better alignment.
@@ 6.2 Principal Component Analysis
We further utilize the dimensionality reduction technique to visualize the intermediate layer latent outputs of the model across different languages. PCA (Principal Component Analysis) (Pearson, 1901) can be used in some scenarios where logit lens doesn’t work. Principal components are a few linear combinations of the original variables that maximally explain the variance of all the variables (Greenacre et al., 2022). We utilize PCA to map the latent logits into the two-dimensional representation. Based on the patterns shown in Figure 1, we report layer 20, 25, 30 and the last layer as four representative layers in Figure 2. We have the following findings: (1) The points of different languages follow the similar patterns in layer 20 and layer 25, where English latent outputs have appeared and outputs in the target language haven’t appeared. We further calculate the Pearson correlation coefficient of 1 dimension PCA results (Appendix C). There is a strong linear correlation between representations of different languages, which also indicates an uniform latent representation pattern during inference process; (2) Representations belong to different languages exhibit greater distance from each other after training; (3) The results of the last layer is similar because of the same possible outputs; (4) Based on Pearson coefficient reported in Appendix C, the correlation between low-resource languages (hi, th, sw, ms) and other high-resource languages significantly improves, which suggests better alignment with English.